OpenStudy (anonymous):

(7/6t^2)- (1/3t) how do I subtract this?

OpenStudy (anonymous):

\[ \frac{7 t^2}{6}-\frac{t}{3}=\frac{1}{6} \left(7 t^2-2 t\right) \]

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

That is it

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

Is this (1/3t) (1/3) t or 1/(3 t)?